# A predictive modular approach to constraint satisfaction under uncertainty - with application to glycosylation in continuous monoclonal antibody biosimilar production

### v1

@@ Abstract
The paper proposes a modular-based approach to constraint handling in process optimization and control. This is partly motivated by the recent interest in learning-based methods, e.g., within bioproduction, for which constraint handling under uncertainty is a challenge. The proposed constraint handler, called predictive filter, is combined with an adaptive constraint margin and a constraint violation cost monitor to minimize the cost of violating soft constraints due to model uncertainty and disturbances. The module can be combined with any controller and is based on minimally modifying the controller output, in a least squares sense, such that constraints are satisfied within the considered horizon. The proposed method is computationally efficient and suitable for real-time applications. The effectiveness of the method is illustrated through a realistic simulation case study of glycosylation constraint satisfaction in continuous monoclonal antibody biosimilar production using Chinese hamster ovary cells, for which the metabolic network model consists of 23 extracellular metabolites and 126 reactions.
@@
Constraint handling , predictive filter , adaptive constraint margin , learning , feedback control , glycosylation , monoclonal antibody , crucial quality attribute , process analytical technology
@@ 1 Introduction
Recently, learning-based optimization and control methods have received significant interest [3], mainly due to their strong learning ability in highly uncertain systems. Such methods have been developed and applied in applications ranging from robotics to biochemical processes. Some examples include imitation learning [49], learning-based model predictive control [26], reinforcement learning [38, 50, 29, 47], Bayesian optimization [54, 40], iterative learning control and optimization [69, 1], etc. Learning based algorithms are in most cases explorative and hence prone to violate constraints that are imposed e.g., for safety or economical reasons.
@@
In process systems, constraints are typically imposed on variables such as concentrations, temperatures, pressures, etc., and it is often optimal to operate at or close to such constraints. The constraints can in most cases be temporally violated, but then often with significant economic costs. Due to large uncertainties in many process systems, such high-cost constraints are likely to be violated, especially when using explorative learning algorithms.
@@
To minimize the cost of constraint violations in the presence of uncertainty, e.g., disturbances, we propose in this work to combine a predictive filter [64] with an adaptive constraint margin and a constraint violation cost monitor. Predictive filters are used, for example, in autonomous driving applications [59]. The margin adaptation employed in this work is based on the margin optimization proposed for economic model predictive control in [61]. Our main contribution is the combination of the predictive filter and the margin adaptation such that the modular constraint handler can minimize the cost of soft constraint violations by learning the optimal constraint margins directly from data. In order to detect any change in uncertainties that impact on constraint violations, we propose to employ a constraint violation cost monitor and adapt the constraint margin so as to minimize the cost.
@@
A particular strength of the proposed method is that it provides a simple and universal module that can be directly applied to systems with arbitrary controllers or optimizers. Notably, the proposed method only alters the control signal when necessary, ensuring a minimal modification of the input computed by the controller or optimizer while satisfying constraints, requiring no prior knowledge of the uncertainty. This is of particular importance when employing methods that are explorative, e.g., when learning-based control or optimization methods are applied. The proposed method is computationally efficient and therefore suitable for real time applications.
@@
There exist a number of methods to handle constraints in the presence of uncertainty. In standard Model Predictive Control (MPC), e.g., [10], a fixed nominal model is employed for constrained optimization and uncertainty is dealt with implicitly by adding output feedback. If some prior knowledge on the uncertainty is available, e.g., that it belongs to a given stochastic distribution or some bounded set, stochastic or robust MPC can be employed [6, 5, 41, 70, 45, 36, 7, 42]. When an a priori model of the system dynamics is lacking, model-free methods that give priority to safety and robustness can be applied, such as safe reinforcement learning [16, 22]. A typical case in process control is that a nominal dynamic model is available, but the uncertainty is not well characterized and this causes problems for methods that do not adapt to compensate for observed uncertainty during runtime, such as MPC. In such cases one can improve constraint handling by employing learning and adaptive based methods [21, 12, 61, 8, 43, 13]. However, essentially all such methods rely on particular controller structures, such as different variants of MPC, and are therefore not applicable to other control algorithms. Predictive filters [59, 64], on the other hand, are modular and can be combined with essentially any control algorithm but is focused on satisfying hard safety constraints. Predictive filters have recently been considered also for dealing with soft constraints, but then primarily to avoid infeasibilities [65, 62]. There also exist works on robust extensions of the predictive safety filter, e.g., [66], but these are all based on prior knowledge of the uncertainty. We note that constraints also can be handled with other methods, such as control barrier functions [4] and Hamilton–Jacobi reachability analysis [20]. The interested reader is referred to [63] and [11] for an overview of constraint handling methods. Finally, in the context of learning-based control, there exist works focusing on parameter adaptation with economic considerations [29, 31, 23, 30, 2, 71, 46]. However, rather than high-cost constraint satisfaction for arbitrary systems, they consider obtaining operational conditions to optimize economical performance using learning based methods. To our knowledge there exist no methods for hard and soft constraint satisfaction that can be applied to systems with arbitrary controllers while adapting optimal constraint margins for the soft constraints directly from data, requiring no prior knowledge on the uncertainty.
@@
Preliminary results have been presented in [68] using examples involving water tank temperature control and fed-batch penicillin fermentation optimization. Here we extend these results, present a realistic case study of glycosylation constraint satisfaction in continuous monoclonal antibody biosimilar production based on a realistic metabolic network and glycosylation models and also demonstrate the advantages of feedback control of glycosylation under uncertainty.
@@
The rest of the paper is organized as follows. We first provide an introductory example of glycosylation constraint satisfaction in Section 2. We formulate the problem considered in Section 3, and introduce the proposed constraint handler in Section 4. In Section 5, we illustrate the effectiveness by revisiting the introductory example. Finally, we provide conclusions and discussion of the work in Section 6.
@@ 2 Introductory example
We present an introductory example to illustrate the impact of uncertainty on constraint handling and the need for constraint handling methods that can learn optimal constraint margins from real time data and, furthermore, be used in combination with arbitrary controllers.
@@
Glycosylation is a common method to modify the properties of proteins by adding carbohydrate components to form glycoproteins. We here consider glycosylation in continuous therapeutic monoclonal antibody (mAb) biosimilar production and in particular constraints on the level of $G0$ glycoforms. Regulatory agencies impose a required range that the $G0$ level should be within for safety and efficacy reasons [15, 33, 19]. Assume process optimization gives an optimal $G0$ percentage level $P_{G0,opt}=35\%$, which coincides with the lower limit of the $G0$ percentage. If the $G0$ percentage $P_{G0}$ is less than the optimal value $P_{G0,opt}$, it will induce a significant economic loss. To track $P_{G0,opt}$ in the presence of disturbances, we apply a simple feedback control strategy to adjust glucose concentration in the feeding medium based on in-line measurements of $P_{G0}$. More details will be provided when we return to this process towards the end of the paper.
@@
As shown in Figure 1 (in blue), when applying the feedback controller to track $P_{G0,opt}$, there are as expected frequent violations of the constraint $P_{G0}>P_{G0,opt}$ that leads to a significant economic loss as can be seen in the lower part of the figure. The main reason for the repeated constraint violations is the presence of various disturbances.
@@
In order to handle the constraints, we add a separate control module named predictive filter [64] which modifies the control signal generated by the feedback controller if a constraint violation is predicted within the considered horizon. The prediction requires a dynamic model that can compute future values of the $P_{G0}$ given the current state of the system. For this purpose, we here employ the model established in [72, 73] and [25], and the constraint is imposed on the glucose consumption rate $q_{Glc}$ which is assumed proportional to the glucose consumption $P_{G0}$ [72]. As can be seen from Figure 1 (in red), the constraint violations are now somewhat less severe but still frequent, mainly due to the fact that future disturbances can not be predicted combined with model uncertainty, i.e., discrepancies between the model used for predictions and the true process.
@@
Since constraint violations are inevitable in the presence of unmeasured disturbances and model uncertainty, it is advisable to introduce a constraint margin, i.e., to force nominal operation some distance away from the constraint. For this purpose, we introduce a constraint margin $\theta\in\mathbb{R}$ such that the modified constraint in the predictive filter is $q_{Glc}\geq q_{Glc,ref}+\theta$. As shown in Figure 1 (middle, in yellow), with $\theta=0.1$, there is now no violation of $P_{G_{0},opt}$. However, the constraint margin is conservative since the resulting $P_{G_{0}}$ is far from the optimum $P_{G_{0},opt}$, which also results in economic losses. Thus, it is important to adapt the margin such that the combined cost of constraint violations and cost of operating away from the optimum is minimized.
@@ 3 Problem Formulation
where $k\in\mathbb{N}$ represents the time step of the discrete-time system; $x\in\mathbb{R}^{n_{x}}$ is the state vector; $u\in\mathbb{R}^{n_{u}}$ is the control input; $f(\cdot):\mathbb{R}^{n_{u}}\times\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ represents the nonlinear dynamics of the plant; $\Delta_{k}\in\mathbb{R}^{n_{x}}$ represents the uncertainty at time step $k$. A key assumption of this paper is that the magnitude and characteristics of $\Delta_{k}$ is unknown a priori. The system is subject to constraints
@@
Since we aim at solving constraint violation problems in the general case we do not assume linear dynamics, i.i.d. disturbance, etc. Rather we make the following reasonable assumptions
@@ 4 Learning-based predictive filter
In the absence of uncertainty, the constraint satisfaction problem can be solved using a safety predictive filter [59, 64]. In the presence of uncertainty, one can employ robust extensions of the safety predictive filter provided a priori knowledge on the uncertainty magnitude and characteristics are available [66]. However, such a priori knowledge on the uncertainty is usually not available in practice, which leads to challenges when implementing robust constraint handling methods [11].
We here propose a learning-based predictive filter, as illustrated in Figure 2. Basically, we introduce a constraint margin $\theta\in\mathbb{R}^{n_{\lambda}}$ for the high-cost constraints, i.e., $g_{e}(x_{k},u_{k})+\theta\leq 0$, which is learned online directly from available data using the constraint adaptor seeking the margin $\theta$ that minimizes the cost (3). Given a learned $\theta$, the predictive filter then computes the minimum modification of the control input to satisfy the constraint with margin $\theta$. After $\theta$ has converged to an optimal value using the constraint adaptor, a violation cost monitor constantly monitors the high-cost constraint violation to determine whether $\theta$ should be re-learned based on changes in the uncertainty size. Note that although one can also introduce some margins on e.g., optimal operational condition $P_{G_{0},opt}$ in the introductory example in Section 2, the predictive filter provides a more robust and less conservative margin due to its predictive ability.
Figure 2: The learning-based predictive filter framework
We next introduce the different parts of the proposed learning based predictive filter; the predictive filter, the constraint adaptor and the constraint violation cost monitor.
@@ 4.1 Predictive filter
Given a constraint margin $\theta$, we apply a predictive filter to improve constraint satisfaction $g_{e}(y_{k},u_{k})+\theta\leq 0$ by modifying the control signals generated by the controller so as to avoid predicted constraint violations. The filter works in a receding horizon fashion and a new predictions is made at every sample instance.
@@
where $u^{d}_{k}$ is the desired control at the current time step $k$ generated by the controller; $N$ is the prediction horizon of the predictive filter; the subscript $i|k$ represents the $i^{th}$ step ahead prediction with the predictive filter initialized at time step $k$; $V_{p}(\epsilon_{0|k},\epsilon_{1|k},\dots,\epsilon_{N|k})$ represents the penalty function on slack variables $\epsilon_{i|k}\in\mathbb{R}^{n_{\lambda}}$, whose elements are non-negative and represent slack variables on the soft inequality constraints to solve potential infeasibility issues. The predictive filter generates the optimal control sequence $\{u_{i|k}^{*}\}$ of which the first element $u_{0|k}^{*}$ is applied to the plant. The filter is re-applied at the next time step in a receding horizon fashion.
@@ 4.2 Constraint adaptor
The purpose of the constraint adaptor is to learn the optimal constraint margin $\theta$ that minimizes (4) based on online measurements. To obtain gradient information from available, possibly noisy and transient, data we propose to employ finite-difference stochastic approximation (FDSA) [56, 57, 34]. Below we briefly outline the FDSA-based adaptation method. We refer interested readers to [61] for further details.
@@
Note that the FDSA-based method does not require gradient information of $J$ to update $\theta$, but instead adapts $\theta$ directly from online measurements. If
@@
As we will demonstrate later, the FDSA based gradient descent can converge relatively fast to the optimal $\theta$, and hence the optimal margin can be learned efficiently in online applications. Again, we do not provide any stability proof here, but note that the FDSA method will slow down over time which implies that a separation of time scales will appear over time and this will improve stability [61].

### v3

@@ Abstract
The paper proposes a modular-based approach to constraint handling in process optimization and control. This is partly motivated by the recent interest in learning-based methods, e.g., within bioproduction, for which constraint handling under uncertainty is a challenge. The proposed constraint handler, called predictive filter, is combined with an adaptive constraint margin and a constraint violation cost monitor to minimize the cost of violating soft constraints due to model uncertainty and disturbances. The module can be combined with any controller and is based on minimally modifying the controller output, in a least squares sense, such that constraints are satisfied within the considered horizon. The proposed method is computationally efficient and suitable for real-time applications. The effectiveness of the method is illustrated through a realistic case study of glycosylation constraint satisfaction in continuous monoclonal antibody biosimilar production using Chinese hamster ovary cells, employing a metabolic network model consisting of 23 extracellular metabolites and 126 reactions. In the case study, the average constraint-violation cost is reduced by more than 60% compared to the case without the proposed constraint-handling method.
@@
Constraint handling , predictive filter , adaptive constraint margin , learning , glycosylation , perfusion
journal: Journal of Process Control
@@ 1 Introduction
Recently, learning-based optimization and control methods have received significant interest [3], mainly due to their strong learning ability in highly uncertain systems. Such methods have been developed and applied in applications ranging from robotics to biochemical processes. Some examples include imitation learning [49], learning-based model predictive control [26], reinforcement learning [38, 50, 29, 47], Bayesian optimization [54, 40], iterative learning control and optimization [68, 1], etc. Learning-based algorithms are in most cases explorative and hence prone to violate constraints that are imposed e.g., for safety or economical reasons.
@@
In process systems, constraints are typically imposed on variables such as concentrations, temperatures, and pressures, and it is often optimal to operate at or close to such constraints. The constraints can in most cases be temporally violated, but then often with significant economic costs. Due to large uncertainties in many process systems, such high-cost constraints are likely to be violated, especially when using explorative learning algorithms.
@@
To minimize the cost of constraint violations in the presence of uncertainty, e.g., disturbances, we propose in this work to combine a predictive filter [63] with an adaptive constraint margin and a constraint violation cost monitor. Predictive filters are used, for example, in autonomous driving applications [59]. The margin adaptation employed in this work is based on the margin optimization proposed for economic model predictive control in [61]. Our main contribution is the combination of the predictive filter and the margin adaptation such that the modular constraint handler can minimize the cost of soft constraint violations by learning the optimal constraint margins directly from real-time data. In order to detect any change in uncertainties that impact on constraint violations, we propose to employ a constraint violation cost monitor and adapt the constraint margin so as to minimize the cost.
@@
A particular strength of the proposed method is that it provides a simple and universal module that can be directly applied to systems with arbitrary controllers or optimizers. Notably, the proposed method only alters the control signal when necessary, ensuring a minimal modification of the input computed by the controller or optimizer while satisfying constraints, requiring no prior knowledge of the uncertainty. This is of particular importance when employing methods that are explorative, e.g., when learning-based control or optimization methods are applied. The proposed method is computationally efficient and therefore suitable for real-time applications.
@@
There exist a number of methods to handle constraints in the presence of uncertainty. In standard Model Predictive Control (MPC), e.g., [10], a fixed nominal model is employed for constrained optimization and uncertainty is dealt with implicitly by adding output feedback. If some prior knowledge on the uncertainty is available, e.g., that it belongs to a given stochastic distribution or some bounded set, stochastic or robust MPC can be employed [6, 5, 42, 69, 45, 36, 7, 41]. When an a priori model of the system dynamics is lacking, model-free methods that give priority to safety and robustness can be applied, such as safe reinforcement learning [16, 22]. A typical case in process control is that a nominal dynamic model is available, but the uncertainty is not well characterized and this causes problems for methods that do not adapt to compensate for observed uncertainty during runtime, such as MPC. In such cases one can improve constraint handling by employing learning and adaptive based methods [21, 12, 61, 8, 43, 13]. However, essentially all such methods rely on particular controller structures, such as different variants of MPC, and are therefore not applicable to other control algorithms. Predictive filters [59, 63], on the other hand, are modular and can be combined with essentially any control algorithm but are focused on satisfying hard safety constraints. Predictive filters have recently been considered also for dealing with soft constraints, but then primarily to avoid infeasibilities [64, 65]. There also exist works on robust extensions of the predictive safety filter, e.g., [66], but these are all based on prior knowledge of the uncertainty. We note that constraints also can be handled with other methods, such as control barrier functions [4] and Hamilton–Jacobi reachability analysis [20]. The interested reader is referred to [62] and [11] for an overview of constraint handling methods. Finally, in the context of learning-based control, there exist works focusing on parameter adaptation with economic considerations [29, 31, 23, 30, 2, 70, 46]. However, rather than high-cost constraint satisfaction for arbitrary systems, they consider obtaining operational conditions to optimize economical performance using learning based methods. To our knowledge there exist no methods for hard and soft constraint satisfaction that can be applied to systems with arbitrary controllers while adapting optimal constraint margins for the soft constraints directly from data, requiring no prior knowledge on the uncertainty.
We stress that the proposed method is not intended to replace MPC, but rather to serve as a modular constraint-handling layer that can be added on top of any controller. This is especially valuable in settings where constraints must be handled even though the underlying controller—such as legacy industrial logic or modern learning-based strategies—cannot handle constraints explicitly.
@@
Preliminary results have been presented in [67] using examples involving water tank temperature control and fed-batch penicillin fermentation optimization. Here we extend these results, present a realistic case study of glycosylation constraint satisfaction in continuous monoclonal antibody biosimilar production based on a realistic metabolic network and glycosylation models.
@@
The rest of the paper is organized as follows. We first provide an introductory example of glycosylation constraint satisfaction in Section 2. We formulate the considered problem in Section 3, and introduce the proposed constraint handler in Section 4. In Section 5, we illustrate the effectiveness by revisiting the introductory example. Finally, we provide conclusions and discussion of the work in Section 6.
@@ 2 Introductory example
We present an introductory example to illustrate the impact of uncertainty on constraint handling and the need for constraint handling methods that can learn optimal constraint margins from real-time data and, furthermore, be used in combination with arbitrary controllers.
@@
Glycosylation is a common method to modify the properties of proteins by adding carbohydrate components to form glycoproteins. We here consider glycosylation in continuous therapeutic monoclonal antibody (mAb) biosimilar production operating around steady-state conditions and in particular constraints on the level of $G0$ glycoforms. Regulatory agencies impose a required range that the $G0$ level should be within for safety and efficacy reasons [15, 33, 19]. Assume process optimization gives an optimal $G0$ percentage level $P_{G0,opt}=35\%$, which coincides with the lower limit of the $G0$ percentage. If the $G0$ percentage $P_{G0}$ is less than the optimal value $P_{G0,opt}$, the product is substandard and must be rejected, resulting in a significant economic loss. To track $P_{G0,opt}$ in the presence of disturbances, we apply a simple feedback control strategy to adjust glucose concentration in the feeding medium based on real-time estimates of steady-state $P_{G0}$ obtained from in-line measurements of glucose consumption rates. More details will be provided when we return to this process towards the end of the paper.
@@
As shown in Figure 1 (in blue), when applying the feedback controller to track $P_{G0,opt}$, there are, as expected, frequent violations of the constraint $P_{G0}>P_{G0,opt}$, and are reflected in a higher economic loss in the lower part of the figure [^2] Note that we here consider constraints on the instantaneous production in a continuous process, rather than on the batch average as is often considered in biopharmaceutical batch production.. The main reason for the repeated constraint violations is the presence of various disturbances.
@@
In order to handle the constraints, we add a separate control module named predictive filter [63] which modifies the control signal generated by the feedback controller if a constraint violation is predicted within the considered horizon. The prediction requires a dynamic model that can compute future values of the $P_{G0}$ given the current state of the system. For this purpose, we here employ the model established in [71, 72] and [25], and the constraint is imposed on the glucose consumption rate $q_{Glc}$ which is assumed measured and inversely related to $P_{G0}$ [71]. As can be seen from Figure 1 (in red), the constraint violations are now somewhat less severe but still frequent, mainly due to the fact that future disturbances can not be predicted combined with model uncertainty, i.e., discrepancies between the model used for predictions and the true process.
@@
Since constraint violations are inevitable in the presence of unmeasured disturbances and model uncertainty, it is advisable to introduce a constraint margin, i.e., to force nominal operation some distance away from the constraint. For this purpose, we introduce a constraint margin $\theta\in\mathbb{R}$ such that the modified constraint in the predictive filter is $q_{Glc}\leq q_{Glc,ref}-\theta$. As shown in Figure 1 (middle, in yellow), with $\theta=0.1$, there is now no violation of $P_{G_{0},opt}$. However, the constraint margin is conservative since the resulting $P_{G_{0}}$ is far from the optimum $P_{G_{0},opt}$, which also results in economic losses. Thus, it is important to adapt the margin such that the combined cost of constraint violations and cost of operating away from the optimum is minimized.
@@ 3 Problem Formulation
where $k\in\mathbb{N}$ represents the time step of the discrete-time system; $x\in\mathbb{R}^{n_{x}}$ is the state vector; $u\in\mathbb{R}^{n_{u}}$ is the control input; $f(\cdot):\mathbb{R}^{n_{u}}\times\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ represents the nonlinear dynamics of the process; $\Delta_{k}\in\mathbb{R}^{n_{x}}$ represents the uncertainty at time step $k$. A key assumption of this paper is that the magnitude and characteristics of $\Delta_{k}$ is unknown a priori. The system is subject to constraints
@@
Since we aim at solving constraint violation problems in the general case we do not assume linear dynamics, independent and identically distributed disturbance, etc. Rather we make the following reasonable assumptions
@@ 4 Learning-based predictive filter
In the absence of uncertainty, the constraint satisfaction problem can be solved using a safety predictive filter [59, 63]. In the presence of uncertainty, one can employ robust extensions of the safety predictive filter provided a priori knowledge on the uncertainty magnitude and characteristics are available [66]. However, such a priori knowledge on the uncertainty is usually not available in practice, which leads to challenges when implementing robust constraint handling methods [11]. Robust methods also tend to be overly conservative.
We here propose a learning-based predictive filter, as illustrated in Figure 2. Basically, we introduce a constraint margin $\theta\in\mathbb{R}^{n_{\lambda}}$ for the high-cost constraints, i.e., $g_{e}(x_{k},u_{k})+\theta\leq 0$, which is learned online directly from available data using the constraint adaptor seeking the margin $\theta$ that minimizes the cost (4). Given a learned $\theta$, the predictive filter then computes the minimum modification of the control input to satisfy the constraint with margin $\theta$. After $\theta$ has converged to an optimal value using the constraint adaptor, a violation cost monitor constantly monitors the high-cost constraint violation to determine whether $\theta$ should be re-learned based on changes in the uncertainty size and characteristics.
Figure 2: The learning-based predictive filter framework, where $x$ denotes the measured or estimated states.
We next introduce the different parts of the proposed learning based predictive filter: the predictive filter, the constraint adaptor and the constraint violation cost monitor.
@@ 4.1 Predictive filter
Given a constraint margin $\theta$, we apply a predictive filter to improve constraint satisfaction $g_{e}(y_{k},u_{k})+\theta\leq 0$ by modifying the control signals generated by the controller so as to avoid predicted constraint violations. The filter works in a receding horizon fashion, and a new prediction is made at every sample instance.
@@
where $u^{d}_{k}$ is the desired control at the current time step $k$ generated by the controller; $N$ is the prediction horizon of the predictive filter; the subscript $i|k$ represents the $i^{th}$ step ahead prediction with the predictive filter initialized at time step $k$; $V_{p}(\epsilon_{0|k},\epsilon_{1|k},\dots,\epsilon_{N|k})$ represents the penalty function on slack variables $\epsilon_{i|k}\in\mathbb{R}^{n_{\lambda}}$, whose elements are non-negative and represent slack variables on the soft inequality constraints to solve potential infeasibility issues. The predictive filter generates the optimal control sequence $\{u_{i|k}^{*}\}$ of which the first element $u_{0|k}^{*}$ is applied to the process. The filter is re-applied at the next time step in a receding horizon fashion.
@@ 4.2 Constraint adaptor
The purpose of the constraint adaptor is to learn the optimal constraint margin $\theta$ that minimizes (4) based on online measurements. To obtain gradient information from available, possibly noisy and transient, data, we propose to employ finite-difference stochastic approximation (FDSA) [56, 57, 34]. Below we briefly outline the FDSA-based adaptation method. We refer interested readers to [61] for further details.
@@
The parameters in FDSA can be selected following standard guidelines [56, 57]. Trollberg et al. [61] also discuss practical heuristics for parameter tuning in FDSA implementations. Also note that the FDSA-based method uses a data-based finite-difference approximation that estimates the gradient directly from measurement based evaluations of $J$ and therefore does not require any explicit derivatives of $J$. If
@@
As we will demonstrate below, the FDSA-based gradient descent can converge relatively fast to the optimal $\theta$, and hence the optimal margin can be learned efficiently in online applications. Again, we do not provide any stability proof here, but note that the FDSA method will slow down over time which implies that a separation of time scales will appear over time and this will improve stability [61].